OpenStudy (anonymous):

Which points are on the graph of the function rule f(x) = 10 – 4x?

jimthompson5910 (jim_thompson5910):

Are you given a list of choices?

OpenStudy (anonymous):

18, –2), (10, 0), (2, 2) (–18, 2), (–10, 0), (–2, –2) (–2, 18), (0, 10), (2, 2) (2, –18), (0, –10), (–2, –2)

jimthompson5910 (jim_thompson5910):

f(x) = 10 – 4x is the same as y = 10 - 4x

jimthompson5910 (jim_thompson5910):

Let's check and see if the point (-2,-2) is on the graph of y = 10 - 4x y = 10 - 4x y = 10 - 4(-2) .. plug in x = -2 -2 = 10 - 4(-2) ... plug in y = -2 -2 = 10 + 8 -2 = 18 ... this is false so (-2,-2) is NOT on y = 10 - 4x

jimthompson5910 (jim_thompson5910):

Do you see how I worked that out?